# Cost and accuracy of long-term graph memory in distributed LLM-based multi-agent systems

### v1

@@ 4.3 Comparative evaluation
Applying the statistical pareto efficiency framework to our experimental data allows us to answer the research questions.
Which memory framework provides the best balance between knowledge retention as well as computational and financial cost in MAS?
Results confirm that mem0 offers the better balance for MAS. While Graphiti showed a raw accuracy increase of 3.6% in the unconstrained experiment and 2% in the constrained experiment, the statistical insignificance renders this advantage negligible. Conversely, the financial cost differential is deterministic and substantial: Graphiti imposed an increased financial burden of 40.2% during the unconstrained experiment. Thus, mem0 is established as the Pareto-optimal solution.
How do knowledge retention as well as computational and financial cost vary when MAS operate in a hybrid cloud–edge environment?
@@ 4.4 Limitations and future work
Consistent with the recommendation for researchers to explore dynamic environments, it is necessary to contextualize the boundaries of the current implementation. While this research offers a first contribution to the nascent field of DMAS memory, specific limitations suggest critical avenues for future work.
@@
Secondly, the chosen network constraints had limited impact on the overall DMAS performance. Future experiments could introduce a peer-to-peer DMAS scenario where agents must decide whether to collaborate or solve tasks locally based on dynamic monitoring of peer disconnections, bandwidth, latency, and jitter.
Thirdly, the DMAS responded to a significant number of questions with IDK. Future work should investigate the root causes of these responses. For example, running experiments with a standard RAG system based on a vector database, for example Qdrant, as a control group could clarify how mem0’s compression algorithm or Graphiti’s knowledge graph engine influences result quality. In that regard, a second conversation could be processed to compare cost and accuracy across different sets of memory.
@@ A.2 Prompts
A.3 Example of retrieved memories in the responder agent

### v2

@@ 4.3 Comparative evaluation
Applying the statistical pareto efficiency framework to our experimental data answered the research questions.
Which memory framework provides the best balance between knowledge retention as well as computational and financial cost in MASs?
Results confirmed that mem0 offers the better balance for MAS. While Graphiti showed a raw accuracy increase of 3.6% in the unconstrained experiment and 2% in the constrained experiment, the statistical insignificance renders this advantage negligible. Conversely, the financial cost differential is deterministic and substantial: Graphiti imposed an increased financial burden of 40.2% during the unconstrained experiment. Thus, mem0 is established as the Pareto-optimal solution.
How do knowledge retention as well as computational and financial cost vary when MASs operate in a hybrid cloud–edge environment?
@@ 4.4 Limitations and future work
While this research offers a first contribution to the nascent field of DMAS memory, specific limitations suggest critical avenues for future work.
@@
Secondly, the chosen network constraints had limited impact on the overall DMAS performance. Future experiments could introduce a peer-to-peer DMAS scenario where agents must decide in response to their environment whether to individually or collaboratively solve tasks using dynamic monitoring of peer disconnections, bandwidth, latency, and jitter.
Thirdly, the DMAS responded to a significant number of questions with IDK. Future work should investigate the root causes of these responses. For example, running experiments with a standard RAG system based on a vector database such as Qdrant as a control group could clarify how mem0’s compression algorithm or Graphiti’s knowledge graph engine influences result quality. In that regard, a second conversation could be processed to compare cost and accuracy across different sets of memory.
@@ A.2 Prompts
A.3 Example of memories retrieved from mem0 and used by the responder agent